The following reaction 2H_2S(g) 2H_2(g) + S_2(g),
Answer #1

------------------ 2H2S(g) <======> 2H2(g) + S2(g) : Kc = 1.67x10-7  

Initl.conc(M): 0.100 ------------------- 0.100, ----- 0.00

change (M): - 2y, --------------------- +2y, -------- +y

eqm.conc(M): (0.100 - 2y), --------- (0.100+2y), y

Kc = 1.67x10-7 = [H2(g)]2x[S2(g)] / [H2S(g)]2

=>  1.67x10-7 = (0.100+2y)2 x y / (0.100 - 2y)2

=> 4y3 + 0.4y2 + 0.01y - 1.67x10-9 = 0

On solving the above equation we get

y = 1.67x10-7 M

Hence equilibrium concentration ofS2(g), [S2(g) = y = 1.67x10-7 M (answer)